OpenStudy (anonymous):

A gardener has 200 m of fencing to enclose two adjacent rectangular plots. What dimensions will produce a maximum enclosed area?

OpenStudy (anonymous):

1. Draw the picture. 2. Formulate the equation.\[A(x)=x(\frac{200-3x}{2})=-\frac{3}{2}x^2+100x\] 3. Note that the function is a parabola opening down, so to find the x value that gives the maximum area, find the x value of the vertex.

OpenStudy (anonymous):

the formula of the vertex is a(x-h)2 + k? Am i right? @AnimalAin

OpenStudy (anonymous):

The x is 75 :)

OpenStudy (anonymous):

Right? @AnimalAin

OpenStudy (anonymous):

I get 100/3, with the width 50.

OpenStudy (anonymous):

OK, I had a little bit of a computer problem here.... let me try again.

OpenStudy (anonymous):

Your formula is related to the vertex, but we need to know the maximum area to make it work. It is probably possible, but cumbersome.

OpenStudy (anonymous):

To find the x value of the vertex for y= ax^2 +bx+c, let x=-b/2a. In this case,\[x=\frac{-100}{-3}=\frac{100}{3}\]\[\implies \frac{200-3x}{2}=50\]

OpenStudy (anonymous):

Isn't -b=3/2? and a=100

OpenStudy (anonymous):

Note that a is the coefficient of the x^2 term, and b is the coefficient of the x term.

OpenStudy (anonymous):

Ahhh. :) hmm. Okaay, i forgot. Sorry ^_^v is the x value the last answer? :/

OpenStudy (mathstudent55):

\(-\dfrac{b}{2a} = - \dfrac{100}{2\left( -\frac{3}{2} \right)} = \dfrac{100}{ 3} \)

OpenStudy (anonymous):

I think it asks for the maximum area, which would be A(100/3). You can get it by substitution, or by multiplying the length and width (easier way to get the same answer). I think the answer is about 5000/3 m^2.

OpenStudy (anonymous):

Getting late here; gotta get up in the AM. Hope I helped you out. Do math every day.

OpenStudy (anonymous):

Thank yoou so much for the help. You really helped alot. Sleep well ^_^v God bless.